OpenStudy (anonymous):

Determine ∫(sin^2 x) (cos^5 x) dx.

OpenStudy (anonymous):

The equation should look like this - \[\int\limits (\sin ^{2}x) (\cos ^{5}x) dx\] And the answer is \[\sin ^{3}x /3 - 2\sin ^{5}x/5 + \sin ^{7}x/7+ c\]

OpenStudy (anonymous):

i think the gimmick is you split off a cosine and write \[\int \sin^2(x) \cos^4(x) \cos(x)dx\]

OpenStudy (anonymous):

then write \[\int \sin^2(x)(1-\sin^2(x))^2 \cos(x)dx\] and now a simple u-sub via \[u=\sin(x)\] etc

OpenStudy (anonymous):

get \[\int u^2(1-u^2)^2du\] and should be easy from there, yes?

OpenStudy (anonymous):

then the answer is : (u^3 + u / 3) + C